# Learning with Calibration: Exploring Test-Time Computing of Spatio-Temporal Forecasting

### v1

@@ 5.4 Flexibility Study (RQ3)
STKEC $25.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.05}}$ $40.11{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.13}}$ $33.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.64}}$ $16.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.04}}$ $26.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.07}}$ $22.33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.16}}$ $5.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.15}}$ $5.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.10}}$ $52.40{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.10}}$
@@
$24.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $39.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.05}}$ $31.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.04}}$ $16.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ $26.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ $21.88{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.06}}$ $3.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.09}}$ $4.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.08}}$ $43.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.90}}$
$\Delta$downarrow 4.6% $\downarrow$ 2.7% $\downarrow$ 4.7% $\downarrow$ 1.2% $\downarrow$ 1.3% $\downarrow$ 2.0% $\downarrow$ 29.2% $\downarrow$ 25.2% $\downarrow$ 17.5%
@@ 5.5 Mechanism & Robustness Study (RQ4)
Figure 7: Left: Strategy comparison. Middle: Effect of $lr$ w.r.t $m$ Right: Effect of $n$ w.r.t $s$.
@@
Strategy Study. We compare different strategies: 1) simple nonlinear time domain calibration (Time), 2) learning only phase or amplitude modulation factors (Pha. / Amp.), 3) node-share modeling (Node), and (4) random selection or retrieval of the most similar samples (Rand. / Sim.). As shown in Figure 7 left, we observe: ❶ Frequency-domain calibration significantly outperforms time-domain calibration, with amplitude modulation being the primary contributor; ❷ Sharing nodes leads to performance degradation due to spatial heterogeneity in spatio-temporal data; ❸ Random sample selection reduces performance, and retrieving similar samples offers negligible gains while incurring higher computational cost. Our proposed update strategy is already near-optimal.
@@ 6 Conclusion
Supplementary material
Learning with Calibration: Exploring
Test-Time Computing of Spatio-Temporal Forecasting
[…]
@@ Appendix A More Related Work
Many recent forecasting models leverage spectral (Fourier or wavelet) representations to capture periodic or multiscale patterns in spatio-temporal data. For example, PastNet [79] integrates a Fourier-domain convolutional operator to embed physical inductive biases, achieving state-of-the-art results in weather and traffic prediction. FourierGNN [85] builds a learnable Fourier-graph operator that conducts graph convolutions in the frequency domain, reducing convolutional complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$. Wavelet-based methods like WDNO [25] perform diffusion modeling in the wavelet domain to capture abrupt spatio-temporal changes and multi-resolution features. In the pure time-series setting, approaches such as FITS [84] interpolate in the complex Fourier domain and discard negligible high-frequency components to maintain accuracy with very few parameters, and TimeKAN [26] explicitly decomposes multivariate series into multiple frequency bands using Kolmogorov–Arnold networks. These works demonstrate how frequency-domain learning can improve forecasting efficiency and accuracy by isolating dominant spectral components. Different from these methods, our method combines spectral domain feature extraction with calibration-aware test-time computation to achieve reliable and calibrated forecasts even under changing conditions.
@@
Traditional online forecasting methods include adaptive filters like Kalman filters and recursive least squares that update linear models on streaming data. Recently, deep-learning approaches have been proposed to handle nonstationarity in an online fashion. For instance, FSNet [55] implements a complementary “fast and slow” learning system: a fast-adapting component for sudden pattern changes and a slow memory component for repeating trends. OneNet [78] runs two parallel neural forecasters (one modeling temporal dependencies, one modeling cross-variable dependencies) and uses reinforcement learning to dynamically weight their predictions under concept drift. These methods continuously update model parameters or ensemble weights as new data arrive. A recent study [35] pointed out the information leakage problem of previous online time series prediction methods, where the model makes predictions and then evaluates them based on the historical time steps that have been back-propagated for parameter updates. By redefining the setting to focus on predicting unknown future steps and evaluating unobserved data points, they propose a two-stream framework for online prediction, DSOF, which is conceptually similar to previous methods, generating predictions in a coarse-to-fine manner through a teacher-student model. Compared with these methods, we focus on the more difficult spatio-temporal predictions while not requiring complex network architecture design. Instead, we propose a calibration-aware framework that focuses on adjusting predictions online instead of learning predictions.
@@
Recent test-time adaptation techniques can be grouped by their adaptation strategy. Entropy minimization methods adjust a trained model to increase prediction confidence on unlabeled test data. For example, [71] propose TENT, which adapts model parameters by minimizing the entropy of its predictions on each test batch and updating batch-normalization layers online. Feature alignment methods recalibrate feature distributions using test inputs; for instance, adaptive batch-normalization techniques re-estimate BN statistics on the target data to align feature distributions without labels. Self-supervised adaptation uses auxiliary tasks on the test data to refine the model. Test-Time Training [64, 18, 24, 76] converts each test input into a self-supervised learning problem (e.g. predicting image rotations) and updates model parameters before making a prediction. Similarly, SHOT [43] freezes the source classifier and updates the feature extractor on unlabeled target data using pseudo-labeling and information maximization. Each of these paradigms improves generalization under distribution shift without access to target labels. In contrast, unlike these methods that exploit self-supervisory information, our spatio-temporal prediction setting can use labels from historical test information, enabling explicit optimization of the objective at test time, ensuring real-time adaptivity.
@@ Appendix B Theoretical Analysis
B.1 Output Perturbation Bound
Theorem 1 (Output Perturbation Bound).
@@ B.1 Output Perturbation Bound
By Parseval’s theorem [53], $\|y^{\prime}-y\|_{2}=\|\Delta Y\|_{2}$, hence:
@@
This theorem guarantees that the calibration-induced perturbation is linearly bounded by the modulation parameters $\epsilon_{\alpha},\epsilon_{\phi}$. By constraining these parameters (e.g., via regularization during test-time adaptation), SD-Calibrator ensures the calibrated output does not deviate excessively from the original prediction, thereby avoiding overfitting to transient noise. The group-wise parameterization further reduces the effective degrees of freedom (from $\mathcal{O}(NM)$ to $\mathcal{O}(NG)$), inherently limiting the risk of over-parameterization.
@@ B.2 Controlled Descent on Streaming Memory Queues
According to the descent Lemma [48], this implies:
@@ C.1 Spectral Domain Calibrator
0:  Pre-trained backbone $f_{\theta}$, Test input $x$, Horizon length $T$, Number of nodes $N$, Groups $G$
@@
2:  Compute $M\leftarrow\tfrac{T}{2}+1$vartriangleright I: Spatial-aware Decomposition
3:  Apply real-to-complex FFT along time dimension for each node: $Y_{f}\;\leftarrow\;\mathrm{rFFT}(\hat{y})\;\in\;\mathbb{C}^{N\times M}$
4:  Decompose: $A\leftarrow|Y_{f}|,\quad P\leftarrow\angle Y_{f}$vartriangleright II: Group-wise Modulation
@@
Algorithm Pseudo-code. We further present Algorithm 1 in the form of pytorch pseudo code in Algorithm 2 for easy understanding.
Algorithm 2 PyTorch-style pseudocode: SD-Calibrator Class
{minted}
python class SD_Calibrator(nn.Module): """ Spectral Domain Calibrator with Phase-Amplitude Modulation """
def __init__(self, num_nodes, freq_bins, groups=4): """ Args: num_nodes: number of spatial nodes (N) freq_bins: number of frequency bins (M = T // 2 + 1) groups: number of frequency groups (G) """ super().__init__() self.groups = groups self.group_size = freq_bins // groups
# Learnable offsets for amplitude and phase: (G, N, 1) self.lambda_amp = nn.Parameter( torch.zeros(groups, num_nodes, 1) ) self.lambda_phi = nn.Parameter( torch.zeros(groups, num_nodes, 1) )
def forward(self, y_pred): """ Args: y_pred: prediction from backbone, shape (B, 1, N, T) B defaults to 1, because only one sample can be tested Returns: calibrated prediction, shape (B, 1, N, T) """ B, _, N, T = y_pred.shape y = y_pred[:, 0] # (B, N, T)
Yf = torch.fft.rfft(y, dim=-1) # (B, N, M) A = torch.abs(Yf) P = torch.angle(Yf)
Yf_corr = torch.zeros_like(Yf) for g in range(self.groups): start = g * self.group_size if g == self.groups - 1 end = T // 2 + 1 else end = (g + 1) * self.group_size
lam_a = self.lambda_amp[g].unsqueeze(0) # (1, N, 1) lam_p = self.lambda_phi[g].unsqueeze(0)
A_g = A[:, :, start:end] * (1 + lam_a) P_g = P[:, :, start:end] + lam_p
@@
y_time = torch.fft.irfft(Yf_corr, n=T, dim=-1) return y_time.unsqueeze(1) # (B, 1, N, T)
C.2 Lightning Gradient Update
@@
3:     Receive $x_{t}$, compute default prediction $\hat{y}_{t}=f_{\theta}(x_{t})$vartriangleright I: Streaming Memory Queue
@@
Algorithm Workflow. We further present Algorithm 3 in the form of pytorch pseudo code in Algorithm 4 for easy understanding.
Algorithm 4 PyTorch-style pseudocode: Flash Gradient Update Function
{minted}
python def st_ttc_test(self, test_loader, node_num, T, groups): """ Flash Gradient Update with Streaming Memory Queue """ SDC = SD_Calibrator(node_num, T//2+1, groups).to(self.device) optimizer = torch.optim.Adam(SDC.parameters(), lr=1e-4) loss_fn = self._select_criterion() SMQ, preds = Queue(maxsize=T), []
for x, y in test_loader: x, y = x.to(self.device), y.to(self.device) with torch.no_grad(): y_pred = self.model(x) y_corr = SDC(y_pred)
# Use y_corr for inference y_corr = self.scaler.inverse_transform(y_corr) preds.append(y_corr.cpu().detach().numpy())
SMQ.put((x, y)) if SMQ.full(): x_old, y_old = SMQ.get() with torch.no_grad(): y_pred_old = self.model(x_old)
SDC.train() y_corr_old = SDC(y_pred_old) y_corr_old = self.scaler.inverse_transform(y_corr_old)
loss = loss_fn(y_corr_old, y_old) loss.backward() optimizer.step() optimizer.zero_grad() SDC.eval() return preds

### v2

@@ 5.4 Flexibility Study (RQ3)
STKEC $25.44{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.05}}$ $40.11{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.13}}$ $33.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.64}}$ $16.25{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.04}}$ $6.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.07}} 22$$2.33{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.16}}$ $5.41{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.15}}$ $5.72{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.10}}$ $52.40{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.10}}$
@@
$24.26{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$ $39.02{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.05}}$ $31.73{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.04}}$ $16.05{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ $26.39{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.11}}$ $21.88{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.06}}$ $3.83{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.09}}$ $4.28{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.08}}$ $43.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.90}}$
$\Delta$ $\downarrow$ 4.6% $\downarrow$ 2.7% $\downarrow$ 4.7% $\downarrow$ 1.2% $\downarrow$ 1.3% $\downarrow$ 2.0% $\downarrow$ 29.2% $\downarrow$ 25.2% $\downarrow$ 17.5%
@@ 5.5 Mechanism & Robustness Study (RQ4)
Figure 7: Left: Strategy comparison. Middle: Effect of lrlr w.r.t $m$ . Right: Effect of $n$ w.r.t $s$.
@@
Strategy Study. We compare different strategies: 1) simple nonlinear time domain calibration (Time), 2) learning only phase or amplitude modulation factors (Pha. / Amp.), 3) node-share modeling (Node), and 4) random selection or retrieval of the most similar samples (Rand. / Sim.). As shown in Figure 7 left, we observe: ❶ Frequency-domain calibration significantly outperforms time-domain calibration, with amplitude modulation being the primary contributor; ❷ Sharing nodes leads to performance degradation due to spatial heterogeneity in spatio-temporal data; ❸ Random sample selection reduces performance, and retrieving similar samples offers negligible gains while incurring higher computational cost. Our proposed update strategy is already near-optimal.
@@ 6 Conclusion
Acknowledgment
The first author would like to thank all the anonymous reviewers for their valuable comments and high recognition. Although he has almost lost his passion for this field and is ready to leave, he hopes that this study can still bring something new to the ST community.
This work is mainly supported by the National Natural Science Foundation of China (No. 62402414). This work is also supported by the Huawei Co., Ltd (No. TC20241023027), the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0620), Guangzhou Municipal Science and Technology Project (No. 2023A03J0011), the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628), and a grant from State Key Laboratory of Resources and Environmental Information System, and Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007).
Appendix
@@ Appendix A More Related Work
Many recent forecasting models leverage spectral (Fourier or wavelet) representations to capture periodic or multiscale patterns in spatio-temporal data. For example, PastNet (Wu et al., 2024) integrates a Fourier-domain convolutional operator to embed physical inductive biases, achieving state-of-the-art results in weather and traffic prediction. FourierGNN (Yi et al., 2023) builds a learnable Fourier-graph operator that conducts graph convolutions in the frequency domain, reducing convolutional complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$. Wavelet-based methods like WDNO (Hu et al., 2025) perform diffusion modeling in the wavelet domain to capture abrupt spatio-temporal changes and multi-resolution features. In the pure time-series setting, approaches such as FITS (Xu et al., 2024b) interpolate in the complex Fourier domain and discard negligible high-frequency components to maintain accuracy with very few parameters, and TimeKAN (Huang et al., 2025) explicitly decomposes multivariate series into multiple frequency bands using Kolmogorov–Arnold networks. These works demonstrate how frequency-domain learning can improve forecasting efficiency and accuracy by isolating dominant spectral components. Different from these methods, our method combines spectral domain feature extraction with calibration-aware test-time computation to achieve reliable and calibrated forecasts even under changing conditions.
@@
Traditional online forecasting methods include adaptive filters like Kalman filters and recursive least squares that update linear models on streaming data. Recently, deep-learning approaches have been proposed to handle nonstationarity in an online fashion. For instance, FSNet (Pham et al., 2023) implements a complementary “fast and slow” learning system: a fast-adapting component for sudden pattern changes and a slow memory component for repeating trends. OneNet (Wen et al., 2023) runs two parallel neural forecasters (one modeling temporal dependencies, one modeling cross-variable dependencies) and uses reinforcement learning to dynamically weight their predictions under concept drift. These methods continuously update model parameters or ensemble weights as new data arrive. A recent study (Lau et al., 2025) pointed out the information leakage problem of previous online time series prediction methods, where the model makes predictions and then evaluates them based on the historical time steps that have been back-propagated for parameter updates. By redefining the setting to focus on predicting unknown future steps and evaluating unobserved data points, they propose a two-stream framework for online prediction, DSOF, which is conceptually similar to previous methods, generating predictions in a coarse-to-fine manner through a teacher-student model. Compared with these methods, we focus on the more difficult spatio-temporal predictions while not requiring complex network architecture design. Instead, we propose a calibration-aware framework that focuses on adjusting predictions online instead of learning predictions.
@@
Recent test-time adaptation techniques can be grouped by their adaptation strategy. Entropy minimization methods adjust a trained model to increase prediction confidence on unlabeled test data. For example, (Wang et al., 2021) propose TENT, which adapts model parameters by minimizing the entropy of its predictions on each test batch and updating batch-normalization layers online. Feature alignment methods recalibrate feature distributions using test inputs; for instance, adaptive batch-normalization techniques re-estimate BN statistics on the target data to align feature distributions without labels. Self-supervised adaptation uses auxiliary tasks on the test data to refine the model. Test-Time Training (Gandelsman et al., 2022; Hardt and Sun, 2024; Sun et al., 2020; Wang et al., 2025a) converts each test input into a self-supervised learning problem (e.g. predicting image rotations) and updates model parameters before making a prediction. Similarly, SHOT (Liang et al., 2020) freezes the source classifier and updates the feature extractor on unlabeled target data using pseudo-labeling and information maximization. Each of these paradigms improves generalization under distribution shift without access to target labels. In contrast, unlike these methods that exploit self-supervisory information, our spatio-temporal prediction setting can use labels from historical test information, enabling explicit optimization of the objective at test time, ensuring real-time adaptivity.
@@ Appendix B Theoretical Analysis
B.1 Approximate Bound on Output Perturbation
Theorem 1 (Approximate Bound on Output Perturbation).
@@ B.1 Approximate Bound on Output Perturbation
By Parseval’s theorem (Parseval, 1806), $\|y^{\prime}-y\|_{2}=\|\Delta Y\|_{2}$, hence:
@@
This theorem guarantees that the calibration-induced perturbation is sub-linearly bounded by the modulation parameters $\epsilon_{\alpha},\epsilon_{\phi}$. By constraining these parameters (e.g., via regularization during test-time adaptation), SD-Calibrator ensures the calibrated output does not deviate excessively from the original prediction, thereby avoiding overfitting to transient noise. The group-wise parameterization further reduces the effective degrees of freedom (from $\mathcal{O}(NM)$ to $\mathcal{O}(NG)$), inherently limiting the risk of over-parameterization.
@@ B.2 Controlled Descent on Streaming Memory Queues
According to the descent Lemma (Nesterov, 2013), this implies:
@@ C.1 Spectral Domain Calibrator
Algorithm Pseudo-code. We further present Algorithm 1 in the form of pytorch pseudo code in Listing LABEL:pytorch_sdc for easy understanding.
C.2 Lightning Gradient Update
Algorithm Workflow. We further present Algorithm 2 in the form of pytorch pseudo code in Listing LABEL:pytorch_fgu for easy understanding.
@@ C.2 Lightning Gradient Update
0: Pre-trained backbone $f_{\theta}$, Test input $x$, Horizon length $T$, Number of nodes NN, Groups $G$
@@
[…]
3: Apply real‐to‐complex FFT along time dimension for each node: $Y_{f}\;\leftarrow\;\mathrm{rFFT}(\hat{y})\;\in\;\mathbb{C}^{N\times M}$
[…]
@@
Listing 1: PyTorch-style pseudocode: SD-Calibrator Class
class SD_Calibrator($n$.Module):
Spectral Domain Calibrator with Phase-Amplitude Modulation
def __init__(self, num_nodes, freq_bins, groups=4):
Args:
num_nodes: number of spatial nodes (N)
freq_bins: number of frequency bins (M = T // 2 + 1)
groups: number of frequency groups (G)
super().__init__()
self.groups = groups
self.group_size = freq_bins // groups
# Learnable offsets for amplitude and phase: (G, N, 1)
self.lambda_amp = $n$.Parameter(
torch.zeros(groups, num_nodes, 1)
self.lambda_phi = $n$.Parameter(
torch.zeros(groups, num_nodes, 1)
def forward(self, y_pred):
Args:
y_pred: prediction from backbone, shape (B, 1, N, T)
B defaults to 1, because only one sample can be tested
Returns:
calibrated prediction, shape (B, 1, N, T)
B, _, N, T = y_pred.shape
y = y_pred[:, 0] # (B, N, T)
Yf = torch.fft.rfft(y, dim=-1) # (B, N, M)
A = torch.abs(Yf)
P = torch.angle(Yf)
Yf_corr = torch.zeros_like(Yf)
for g in range(self.groups):
start = g * self.group_size
if g == self.groups - 1
end = T // 2 + 1
else
end = (g + 1) * self.group_size
lam_a = self.lambda_amp[g].unsqueeze(0) # (1, N, 1)
lam_p = self.lambda_phi[g].unsqueeze(0)
A_g = A[:, :, start:end] * (1 + lam_a)
P_g = P[:, :, start:end] + lam_p
@@
y_time = torch.fft.irfft(Yf_corr, n=T, dim=-1)
return y_time.unsqueeze(1) # (B, 1, N, T)
@@
[…]
@@
Listing 2: PyTorch-style pseudocode: Flash Gradient Update Function
def st_ttc_test(self, test_loader, node_num, T, groups):
Flash Gradient Update with Streaming Memory Queue
SDC = SD_Calibrator(node_num, T//2+1, groups).to(self.device)
optimizer = torch.optim.Adam(SDC.parameters(), lr=1e-4)
loss_fn = self._select_criterion()
SMQ, preds = Queue(maxsize=T), []
for x, y in test_loader:
x, y = x.to(self.device), y.to(self.device)
with torch.no_grad():
y_pred = self.model(x)
y_corr = SDC(y_pred)
# Use y_corr for inference
y_corr = self.scaler.inverse_transform(y_corr)
preds.append(y_corr.cpu().detach().numpy())
SMQ.put((x, y))
if SMQ.full():
x_old, y_old = SMQ.get()
with torch.no_grad():
y_pred_old = self.model(x_old)
SDC.train()
y_corr_old = SDC(y_pred_old)
y_corr_old = self.scaler.inverse_transform(y_corr_old)
loss = loss_fn(y_corr_old, y_old)
loss.backward()
optimizer.step()
optimizer.zero_grad()
SDC.eval()
return preds